OpenStudy (anonymous):

form a fifth-degree polynomial function with real coefficients such that 5i,1-2i, and -5 are zeros and f(0)=1875

OpenStudy (campbell_st):

well its P(x) = a(x +5)(x -5i)(x + 5i)(x - 1 + 2i)(x - 1 -2i) a complex root has a congugate pair...

OpenStudy (campbell_st):

then after distributing, substitute x = 0 and P(x) = 1875 to find a

OpenStudy (aum):

Roots: 5i,1-2i, and -5 x = 5i square both sides: x^2 = -25 x^2 + 25 = 0 Therefore, (x^2+25) is a factor. x = 1-2i x - 1 = -2i (x-1)^2 = 4 x^2 - 2x + 1 - 4 = 0 x^2 - 2x - 3 = 0 Therefore, x^2 - 2x - 3 is a factor x = -5 Therefore, (x+5) is a factor Multiply (x^2+25), (x^2 - 2x - 3) and (x+5) Once you get the polynomial, write it as a constant a * (the polynomial) f(0)=1875 put x = 0 and f(x) = 1875 to solve for a.

OpenStudy (anonymous):

wait whats thw answer

OpenStudy (aum):

Can't give out answers. I have done most of the work. All you have to do is do a few steps to complete the problem.

OpenStudy (anonymous):

ok so this is the polynomial that I got x^5+3x^4+12x^3+60x^2-325x-375

OpenStudy (anonymous):

what do I do next?

OpenStudy (aum):

okay, your polynomial is P(x) = a * ( x^5+3x^4+12x^3+60x^2-325x-375) put x = 0 and P(x) = 1875 and find 'a'.

OpenStudy (anonymous):

ok

OpenStudy (campbell_st):

I think there is a slight error x = 1 - 2i so x - 1 = -2i squaring both gives \[(x - 1)^2 = 4i^2\] so (x -1)^2 = -4 because i^2 = -1

OpenStudy (aum):

yes, sorry about that. The second factor should be (x^2 - 2x + 5)

OpenStudy (anonymous):

ok

OpenStudy (aum):

So the polynomial is: P(x) = a * ( x^5+3x^4+20x^3+100x^2-125x+625 )

OpenStudy (anonymous):

1875= a * ( 0^5+3(0)^4+20(0)^3+100(0)^2-125(0)+625 )

OpenStudy (aum):

Yes. 1875 = 625 * a a = 1875 / 625 = 3 P(x) = 3(x^5+3x^4+20x^3+100x^2-125x+625)

OpenStudy (anonymous):

ok so thats the answer

OpenStudy (aum):

Yes.

OpenStudy (anonymous):

wait but dont I have to multiply three by everything

OpenStudy (aum):

Either way should be fine. One is in factored form and another is the expanded form but both are the same polynomial.

OpenStudy (anonymous):

THANK YOU SO MUCH

OpenStudy (aum):

You are welcome.